OpenStudy (anonymous):

If two different students are randomly selected from a class of 12 boys and 8 girls, what is the probability that both students are girls?

OpenStudy (anonymous):

what are the answer choices?

OpenStudy (anonymous):

There aren't any answer choices.

OpenStudy (sweetburger):

lel

OpenStudy (anonymous):

Do you have any idea of what the answer is?

OpenStudy (anonymous):

Well, the answer has to be in a common fraction if that helps.

OpenStudy (sweetburger):

number of wanted outcomes ------------------------ number of possible outcomes i meant divide went i put ------

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (raden):

the even of selects both students are girls = 8C2 and number possible outcomes = 20C2 so, P(G,G) = 8C2 / 20C2 = ... just simplify the calculation above

OpenStudy (raden):

use the combination formula nCr = n!/(r!*(n-r)!) so, 8C2 = 8!/(2!6!) = 8.7.6!/(2!6!) = 28 and 20C2 = 20!/(2!18!) = 20.9.18!/(2!18!) = 90 so, P(G,G) = 28/90 = 14/45

OpenStudy (raden):

oppss.. i make a wrong, that should like this 8C2 = 8!/(2!6!) = 8.7.6!/(2!6!) = 28 20C2 = 20!/(2!18!) = 20.19.18!/(2!18!) = 190 so, P(G,G) = 28/190 = 14/95

OpenStudy (raden):

that's 20 * 19 not 20 * 9